### 17 Jan 2024 v4

@@ 3.5.1 Construction of the smoother
The fixed-knot thin plate spline smoother with covariance reweighing is constructed as follows. Denote by $\tilde{\mathbf{R}}\in\mathbb{R}^{n\times m}$ the radial basis matrix with entries:
@@ 3.5.2 Uncertainty quantification
Estimate variances and confidence intervals follow for the thin plate spline estimates through the linearity of the smoother and the assumption of Gaussian noise. Denote by ${\boldsymbol{\Sigma}}\in\mathbb{R}^{n\times n}$ the observation covariance matrix $\operatorname{\mathrm{cov}}(\boldsymbol{y})$ from (21). The covariance matrix of the estimates is given by:
@@ 4.2 Ocean thermal response as a continuous function of time and cross-track distance
Having adjusted the temperature differences for the seasonal change using the model detailed in Section 3.3 and accounted for the variability across profile pairs via Section 3.4, we now apply the thin plate spline smoother to the seasonally adjusted temperature differences to obtain our main scientific result: a characterization of the ocean thermal response to the passage of hurricane-strength tropical cyclones as a continuous function of time and cross-track distance. This function is depicted for data from profile pairs associated with a hurricane-strength tropical cyclone (formally defined as having sustained wind speed of at least 64 knots at $t_{\pi(x_{\mathrm{lon}},x_{\mathrm{lat}})}$, the time of TC passage) at a pressure level of $z=10$ decibars in Figure 7.
The thin plate splines are fit separately at each depth $z$. Within each $z\in\{10,20,\dotsc,200\}$, we select a tuning parameter $\hat{\lambda}^{\text{cv}}_{z}$ guided by leave-one-out cross-validation (LOOCV). Importantly, we do not take the $\hat{\lambda}^{\text{min}}$ that directly minimizes the LOOCV score. Instead, we take the $\lambda<\hat{\lambda}^{\text{min}}$ that produces a LOOCV error slightly (1%) larger than the minimum LOOCV error. This is done because heterogeneity in the thermal response signal causes a direct minimization of the LOOCV score to oversmooth the signal. The full details of the cross-validation procedure, including its mathematical statement, plots illustrating the estimated test error achieved by different $\lambda$, and the chosen $\hat{\lambda}^{\text{cv}}_{z}$, are deferred to Appendix B.
@@
Figure 7: Temperature differences for hurricane-strength profile pairs in the $(d,\tau)$ coordinate system, at a depth of 10 meters, smoothed using thin plate splines. We observe in Figure 7 an unusual positive anomaly at cross-track $d\approx-6$, time difference $\tau\in(3,18)$, due to a set of highly-correlated pairs. After the block covariance correction in Figure 7, the positive anomaly is diminished. A pointwise hypothesis test in Figure 7 completely masks the phenomenon, indicating that it was indeed ocean variability noise unrelated to TCs.
In Figure 7, we plot the surface estimated by applying a thin plate spline smoother without accounting for the variability estimates from Section 3.4, i.e., setting $\mathbf{W}$ to be the identity matrix. We observe a sustained warm anomaly at $d\approx-6$ that is present over several days. Concerningly, this warm anomaly is comparable in magnitude with the expected cooling effect observed at $d\approx 0$.
@@
Fortunately, once we account for the correlation between observations, this warm anomaly essentially disappears in Figure 7. Formally, the correlation adjustment is performed as follows. We obtain an estimate for ${\boldsymbol{\Sigma}}=\operatorname{\mathrm{cov}}(\boldsymbol{y})$ through the Gaussian process model of Section 3.4, and provide this to our covariance-reweighed thin plate spline smoother of Section 3.5 by setting $\mathbf{W}={\boldsymbol{\Sigma}}^{-1}$. Recall that by construction, $\boldsymbol{\Sigma}$} is a block diagonal matrix, with blocks no larger than six entries wide; therefore, ${\boldsymbol{\Sigma}}^{-1}$ may be efficiently computed. For a very small number of temperature differences, the fitted variance is extremely small, on the order of $\leq\exp\{-4.5\}$. We truncate the left tail and filter out these observations to prevent distorting the fit. This omits only 488 observation pairs out of 16025.
@@
Finally, in Figure 7, we perform a pointwise level $\alpha=0.05$ hypothesis test on the thin plate spline fit, using the variance estimates derived in Section 3.5. The issues of performing a pointwise hypothesis test with $N=40000$ points are manifold, and we defer discussion of extensions through which we could make the significance test more sophisticated to Section 12. Nonetheless, we see in Figure 7 a simple fact: that the estimated cooling effect induced by tropical cyclones is of a greater magnitude than the variability of the estimates. Moreover, this cooling effect persists for a sustained period of time, stretching into two weeks following the tropical cyclone passage. We also see a rightward bias in the cooling effect, which may be attributed to the storm’s direction of rotation. [^2] Due to the Coriolis effect, TCs in the Northern Hemisphere will induce a greater cooling effect on the right side of the cross-track angle parameterization than on the left side, due to the combined effects of forward movement and counterclockwise wind rotation. TCs in the Southern Hemisphere, which experience clockwise wind rotation, will induce a greater cooling effect on the left side. In order to display this effect in a consistent manner across hemispheres, we flip the sign of the cross-track angle for TCs in the Southern Hemisphere when producing Figures 612.
@@
An additional phenomenon of scientific interest that may be observed in the thin plate spline fits is that of vertical mixing, in which the warm water close to the ocean surface is forced downwards by the force of the tropical cyclone (Mei and Pasquero, 2013). Although the net change in heat content incidental to the passage of a tropical cyclone is negative, vertical mixing may register as a warming effect at shallower subsurface ocean depths.
@@
Figure 8: Temperature differences for hurricane-strength profile pairs, at depths of $z=10,60,150$, smoothed using thin plate splines. Vertical mixing gives rise to a warming phenomenon on the right side of the tropical cyclone at the intermediate depths. All plots are masked with a pointwise level $\alpha=0.05$ hypothesis test.
@@
The discussion of vertical mixing is further developed in Sections 4.3, 4.4, and 4.5. The full set of covariance-reweighed thin plate spline fits, at all pressure levels $z=10,\dotsc,200$, is deferred to Section 2 of the supplementary material (Hu et al., 2020).
@@ 4.3 Depth-time parameterization
Of additional scientific interest is the characterization of ocean thermal response for a fixed cross-track angle, as a function of time and pressure. To facilitate this analysis, we collect the twenty thin plate spline fits for pressure levels $z=10,\dotsc,200$ from Section 4.2 and “marginalize” them along three sets of cross-track angles: $d\in[-2.5,-1.5]$, $d\in[-0.5,+0.5]$, and $d\in[+1.5,+2.5]$. These three plots are presented in Figure 9.
@@
Figure 9: The covariance-reweighed thin plate spline fits (hurricane-strength TCs) of Section 4.2 are averaged along three sets of cross-track angles $d$ corresponding to the left, center, and right relative to the TC’s forward movement. The resulting estimates are plotted as a function of time since TC passage $\tau$ and pressure level $z$. A pointwise $\alpha=0.05$ hypothesis test is performed and used to mask areas where we fail to reject the null hypothesis.
In each of the marginalized plots, we subset the covariance-reweighed thin plate spline fits from Section 4.2 to each of the cross-track bins and average out the cross-track component. Variance estimates follow immediately (owing to the linearity of averaging) and are used to construct a pointwise $\alpha=0.05$ hypothesis test. We display these marginalized estimates in Figure 9.
@@ 4.4 Depth-cross-track parameterization
We further demonstrate that our continuous-time methodology captures the temporally binned results of Cheng et al. (2015), whose profile pairing process we used in this paper. To introduce temporal binning, we take the covariance-reweighed thin plate spline estimates from Section 4.2 across pressure levels $z=10,\dotsc,200$ and split them into two bins: $\tau\in[0,+3)$ and $\tau\in[+3,+20)$, termed the “forced” stage and “recovery” stage by Cheng et al. (2015). For each of the two bins, we perform the same averaging process as in Section 4.3 to marginalize the estimates over the $\tau$ axis and obtain $\alpha=0.05$ pointwise significance tests.
@@
Figure 10: The covariance-reweighed thin plate spline fits (hurricane-strength TCs) of Section 4.2 are averaged along two sets of time since TC passage $\tau$, and the resulting estimates are plotted as a function of pressure level $z$ and cross-track angle $d$. A pointwise $\alpha=0.05$ hypothesis test is performed and used to mask areas where we fail to reject the null hypothesis.
@@
Several noteworthy phenomena are present in Figure 10. In Figure 10, we observe both the primary cooling effect and the secondary warming effect in the subsurface on the right side of the storm due to vertical mixing, as discussed in Sections 4.2 and 4.3. We also observe a sustained cooling effect in Figure 10, during the period called the “recovery” stage in Cheng et al. (2015). While the primary cooling effect observed in the two time periods is qualitatively consistent with the results presented by Cheng et al. (2015), their results also show a strong warming effect on both sides of the storm during the recovery stage (Figure 10 in Cheng et al., 2015), which is absent from our estimate. Overall, it appears that the recovery stage estimate in Cheng et al. (2015) is “warm biased” in comparison to our estimate. We suspect that this difference is due to our methodology’s more careful accounting of the seasonal mean shift as described in Section 3.3.
4.5 Isosurfaces of the thin plate spline estimates
To provide a holistic understanding of the ocean thermal response structure as a function of depth, cross-track distance, and time since TC passage, we compute isosurface plots of the thin plate spline estimates from Section 3.5. Two sets of isosurfaces are computed; one set for regions of negative temperature differences, and one set for regions of positive temperature differences. These are shaded blue and red, respectively, in Figure 11.
@@ 4.5 Isosurfaces of the thin plate spline estimates
Figure 11: $\pm 0.3^{\circ}\mathrm{C}$ isosurfaces of the thin plate spline fits, where negative temperature differences are shaded in blue and positive temperature differences are shaded in red. Each subplot is labeled with its camera perspective, given in its polar ($\theta$) and azimuthal ($\varphi$) angles.
For each set of isosurfaces, we first mask the thin plate spline estimates that fail to reject a pointwise level $\alpha=0.05$ hypothesis test, using the variance estimates from Section 3.5.2. We then apply the marching cubes algorithm (Lorensen and Cline, 1987; Lewiner et al., 2012; Virtanen et al., 2020) to identify a positive and a negative isosurface of magnitude $0.3^{\circ}\mathrm{C}$. The positive isosurface is colored in red, and the negative in blue.
@@ 4.6 Vertically averaged temperature differences
The seasonally adjusted temperature differences and accompanying thin plate spline fits are presented in Figure 12. In Figure 12, we only show data from pairs associated with hurricane-strength TCs, in contrast to Figure 6, which showed data from all profile pairs regardless of storm intensity. The restriction to hurricane-strength profile pairs here is to provide a fairer comparison to the thin plate spline estimates of Figures 12 and 12, which use the hurricane-strength subset (as in Figure 7).
The thin plate spline fit in Figure 12 conforms with our intuition; we observe a cooling signal, as in Figure 7, but its magnitude is dampened by the vertical averaging procedure. More importantly, Figure 12 does not show any warming effect indicative of vertical mixing. This is in contrast to Figures 7, 9, and 10, which treat depth as a separate dimension of analysis. From an oceanographic and meteorological perspective, this suggests that the vertical mixing-induced warming effect is local in depth; it only appears at some pressure levels, because it is caused by the forcing of warm surface waters downward by the TC. Once we integrate vertically, the local warming effect is cancelled out by a cooling effect near the surface and is no longer visible. From a statistical perspective, this underscores the importance of treating depth as a separate axis of analysis, using the vertical resolution of Argo observations.
@@
Aside from methodological contributions, we have presented several results of scientific interest. In particular, our framework improves upon past scientific literature by modeling the tropical cyclone-induced ocean thermal response as a function of time, pressure, and cross-track angle, in which both the time and cross-track angle components are treated continuously. Through this modeling, we are able to produce fine-grained characterizations of the ocean thermal response at a number of different depths, and we observe differing ocean thermal response as a function of depth. Specifically, we recover the primary effect of surface cooling and secondary effect of subsurface warming due to vertical mixing. By computing isosurfaces on our thin plate spline fits, we are able to characterize these two phenomena as a function of the depth, time, and cross-track distance dimensions simultaneously. Finally, the seasonal mean field component of our methodological framework reveals a nontrivial warming of the ocean near its surface during the TC season, which we estimate and account for in our analysis.
@@
The present methodological framework also admits a number of possible refinements. From a statistical standpoint, one could improve upon our model by considering each temperature profile as a functional observation and modeling the depth component continuously, rather than interpolating to a sequence of depths and fitting separate models at each depth level. This would potentially allow one to also infer changes in mixed layer depth and thermocline stability. Our model could also potentially be improved by replacing the thin plate spline smoother with a spline smoother with a spatially varying smoothing parameter (e.g., Ruppert and Carroll, 2000) or a nonlinear, multiresolution smoother that naturally accounts for heterogeneity in the TC-induced signal (e.g., Daubechies et al., 1999).
@@
An additional extension of statistical and scientific interest involves carefully accounting for float drift. In this paper, we require the baseline and signal profiles to be within $0.2^{\circ}$ of each other and then assume that they share a single location in the analysis, admitting a useful simplification of the fitted covariance function. An extension would involve relaxing the strict closeness requirement of the paired profiles and removing the assumption that the paired profiles are in the same location by more carefully modeling their covariance, taking into account their relative spatial locations. This requires a nontrivial generalization of the Gaussian process model of Kuusela and Stein (2018), but would provide the benefit of increasing the number of available pairs with which to fit the model. Yet another extension would be the consideration of parametric models for the TC wake (Haney et al., 2012). Finally, the fusion of microwave sea surface temperature (mSST; Gentemann et al., 2003; Wentz et al., 2005) observations with Argo profiles might provide an opportunity to produce estimates on the individual-storm level, rather than in terms of the global aggregates, as was done here. mSST observations are very dense in latitude and longitude, but lack the depth penetration provided by the Argo profiles, which are relatively sparse in latitude and longitude. Therefore, mSST and Argo are naturally complementary in the information provided.
@@
Appendix A Derivation of the thin plate spline estimator
In this section, provided for completeness, we motivate and derive the thin plate spline estimator (Duchon, 1977; Wahba, 1980, 1990; Green and Silverman, 1994; Wood, 2003) used in Section 3.5 for completeness. First, we review the natural thin plate spline smoother in Section A.1. Sections A.2 and A.3 then describe versions of the thin plate spline when subject to covariance reweighing (via generalized least squares; Rao, 1973; Draper and Smith, 1998; Nychka, 2000; Ruppert et al., 2003) and when its basis functions are constructed from a fixed knot set (Wahba, 1980; Wood, 2017).
A.1 The natural thin plate spline smoother

### 14 Mar 2024 v5

@@ 3.5.1 Construction of the smoother
The fixed-knot thin-plate spline smoother with covariance reweighting is constructed as follows. Denote by $\tilde{\mathbf{R}}\in\mathbb{R}^{n\times m}$ the radial basis matrix with entries:
@@ 3.5.2 Uncertainty quantification
Estimate variances and confidence intervals follow for the thin-plate spline estimates through the linearity of the smoother and the assumption of Gaussian noise. Denote by ${\boldsymbol{\Sigma}}\in\mathbb{R}^{n\times n}$ the observation covariance matrix $\operatorname{\mathrm{cov}}(\boldsymbol{y})$ from (21). The covariance matrix of the estimates is given by:
@@ 4.2 Ocean thermal response as a continuous function of time and cross-track distance
Having adjusted the temperature differences for the seasonal change using the model detailed in Section 3.3 and accounted for the variability across profile pairs via Section 3.4, we now apply the thin-plate spline smoother to the seasonally adjusted temperature differences to obtain our main scientific result: a characterization of the ocean thermal response to the passage of hurricane-strength tropical cyclones as a continuous function of time and cross-track distance. This function is depicted for data from profile pairs associated with a hurricane-strength tropical cyclone (formally defined as having sustained wind speed of at least 64 knots at $t_{\pi(x_{\mathrm{lon}},x_{\mathrm{lat}})}$, the time of TC passage) at a pressure level of $z=10$ decibars in Figure 7.
The thin-plate splines are fit separately at each depth $z$. Within each $z\in\{10,20,\dotsc,200\}$, we select a tuning parameter $\hat{\lambda}^{\text{cv}}_{z}$ guided by leave-one-out cross-validation (LOOCV). Importantly, we do not take the $\hat{\lambda}^{\text{min}}$ that directly minimizes the LOOCV score. Instead, we take the $\lambda<\hat{\lambda}^{\text{min}}$ that produces a LOOCV error slightly (1%) larger than the minimum LOOCV error. This is done because heterogeneity in the thermal response signal causes a direct minimization of the LOOCV score to oversmooth the signal. The full details of the cross-validation procedure, including its mathematical statement, plots illustrating the estimated test error achieved by different $\lambda$, and the chosen $\hat{\lambda}^{\text{cv}}_{z}$, are deferred to Appendix B.
@@
Figure 7: Temperature differences for hurricane-strength profile pairs in the $(d,\tau)$ coordinate system, at a depth of 10 meters, smoothed using thin-plate splines. We observe in Figure 7 an unusual positive anomaly at cross-track $d\approx-6$, time difference $\tau\in(5,13)$, as well as unusually strong negative differences at time difference $\tau\in(18,20)$. Reweighting by the estimated block covariance in Figure 7 attenuates these anomalies, and a pointwise hypothesis test in Figure 7 masks these phenomena, indicating that they were due to ocean variability noise unrelated to TCs.
@@
The central Figure 7 depicts the thin-plate spline fit, accounting for correlation between observations, at a depth of $z=10$ decibars. Formally, the correlation adjustment is performed as follows. We obtain an estimate for ${\boldsymbol{\Sigma}}=\operatorname{\mathrm{cov}}(\boldsymbol{y})$ through the Gaussian process model of Section 3.4, and provide this to our covariance-reweighted thin-plate spline smoother of Section 3.5 by setting $\mathbf{W}={\boldsymbol{\Sigma}}^{-1}$. Recall that by construction, $\boldsymbol{\Sigma}$} is a block diagonal matrix, with blocks no larger than six entries wide; therefore, ${\boldsymbol{\Sigma}}^{-1}$ may be efficiently computed. For a very small number of temperature differences, the fitted variance is extremely small, on the order of $\leq\exp\{-4.5\}$. We truncate the left tail and filter out these observations to prevent distorting the fit. This omits only 488 observation pairs out of 16025.
@@
In Figure 7, we perform a pointwise level $\alpha=0.05$ hypothesis test on the thin-plate spline fit, using the variance estimates derived in Section 3.5. The issues of performing a pointwise hypothesis test with $N=40000$ points are manifold, and we defer discussion of extensions through which we could make the significance test more sophisticated to Section 12. Nonetheless, we see in Figure 7 a simple fact: that the estimated cooling effect induced by tropical cyclones is of a greater magnitude than the variability of the estimates. Moreover, this cooling effect persists for a sustained period of time, stretching into two weeks following the tropical cyclone passage. We also see a rightward bias in the cooling effect, which may be attributed to the storm’s direction of rotation. [^3] Due to the Coriolis effect, TCs in the Northern Hemisphere will induce a greater cooling effect on the right side of the cross-track angle parameterization than on the left side, due to the combined effects of forward movement and counterclockwise wind rotation. TCs in the Southern Hemisphere, which experience clockwise wind rotation, will induce a greater cooling effect on the left side. In order to display this effect in a consistent manner across hemispheres, we flip the sign of the cross-track angle for TCs in the Southern Hemisphere when producing Figures 612.
To illustrate the importance of estimating the correlations between observations on the final thin-plate spline fit, we plot in Figure 7 the surface estimated by applying a thin-plate spline smoother without accounting for the variability estimates from Section 3.4, i.e., setting $\mathbf{W}$ to be the identity matrix. These estimates are tuned with the same LOOCV procedure. We observe a sustained warm anomaly at $d\approx-6$ that is present over several days. Concerningly, this warm anomaly is comparable in magnitude with the expected cooling effect observed at $d\approx 0$. We also see negative anomalies which occur late in the recovery stage, between $\tau\in(18,20)$. When viewed in context with Figures 7 and 7, we see that adjusting for the correlations reduces the impact of some outlying observations causing these anomalous effects in the naive estimate, and the pointwise hypothesis test masks away the effects almost entirely.
@@
An additional phenomenon of scientific interest that may be observed in the thin-plate spline fits is that of vertical mixing, in which the warm water close to the ocean surface is forced downwards by the force of the tropical cyclone (Mei and Pasquero, 2013). Although the net change in heat content incidental to the passage of a tropical cyclone is negative, vertical mixing may register as a warming effect at shallower subsurface ocean depths.
@@
Figure 8: Temperature differences for hurricane-strength profile pairs, at depths of $z=10,60,150$, smoothed using thin-plate splines. Vertical mixing gives rise to a warming phenomenon on the right side of the tropical cyclone at the intermediate depths. All plots are masked with a pointwise level $\alpha=0.05$ hypothesis test.
@@
The discussion of vertical mixing is further developed in Sections 4.3, 4.4, and 4.5. The full set of covariance-reweighted thin-plate spline fits, at all pressure levels $z=10,\dotsc,200$, is deferred to Section 2 of the supplementary material (Hu et al., 2020).
@@ 4.3 Depth-time parameterization
Of additional scientific interest is the characterization of ocean thermal response for a fixed cross-track angle, as a function of time and pressure. To facilitate this analysis, we collect the twenty thin-plate spline fits for pressure levels $z=10,\dotsc,200$ from Section 4.2 and “marginalize” them along three sets of cross-track angles: $d\in[-2.5,-1.5]$, $d\in[-0.5,+0.5]$, and $d\in[+1.5,+2.5]$. These three plots are presented in Figure 9.
@@
Figure 9: The covariance-reweighted thin-plate spline fits (hurricane-strength TCs) of Section 4.2 are averaged along three sets of cross-track angles $d$ corresponding to the left, center, and right relative to the TC’s forward movement. The resulting estimates are plotted as a function of time since TC passage $\tau$ and pressure level $z$. A pointwise $\alpha=0.05$ hypothesis test is performed and used to mask areas where we fail to reject the null hypothesis.
In each of the marginalized plots, we subset the covariance-reweighted thin-plate spline fits from Section 4.2 to each of the cross-track bins and average out the cross-track component. Variance estimates follow immediately (owing to the linearity of averaging) and are used to construct a pointwise $\alpha=0.05$ hypothesis test. We display these marginalized estimates in Figure 9.
@@ 4.4 Depth-cross-track parameterization
We further demonstrate that our continuous-time methodology captures the temporally binned results of Cheng et al. (2015), whose profile pairing process we used in this paper. To introduce temporal binning, we take the covariance-reweighted thin-plate spline estimates from Section 4.2 across pressure levels $z=10,\dotsc,200$ and split them into two bins: $\tau\in[0,+3)$ and $\tau\in[+3,+20)$, termed the “forced” stage and “recovery” stage by Cheng et al. (2015). For each of the two bins, we perform the same averaging process as in Section 4.3 to marginalize the estimates over the $\tau$ axis and obtain $\alpha=0.05$ pointwise significance tests.
@@
Figure 10: The covariance-reweighted thin-plate spline fits (hurricane-strength TCs) of Section 4.2 are averaged along two sets of time since TC passage $\tau$, and the resulting estimates are plotted as a function of pressure level $z$ and cross-track angle $d$. A pointwise $\alpha=0.05$ hypothesis test is performed and used to mask areas where we fail to reject the null hypothesis.
@@
Several noteworthy phenomena are present in Figure 10. In Figure 10, we observe both the primary cooling effect and the secondary warming effect in the subsurface on the right side of the storm due to vertical mixing, as discussed in Sections 4.2 and 4.3. We also observe a sustained cooling effect in Figure 10, during the period called the “recovery” stage in Cheng et al. (2015). While the primary cooling effect observed in the two time periods is qualitatively consistent with the results presented by Cheng et al. (2015), their results also show a strong warming effect on both sides of the storm during the recovery stage (Figure 10 in Cheng et al., 2015), which is absent from our estimate. Overall, it appears that the recovery stage estimate in Cheng et al. (2015) is “warm biased” in comparison to our estimate. We attribute the difference to our methodology’s accounting of the seasonal mean field and variability estimates, which, when used in the thin-plate spline estimate, attenuate the influence of outlying observations dominated by non-TC oceanographic variability as illustrated in Figure 7.
4.5 Isosurfaces of the thin-plate spline estimates
To provide a holistic understanding of the ocean thermal response structure as a function of depth, cross-track distance, and time since TC passage, we compute isosurface plots of the thin-plate spline estimates from Section 3.5. Two sets of isosurfaces are computed; one set for regions of negative temperature differences, and one set for regions of positive temperature differences. These are shaded blue and red, respectively, in Figure 11.
@@ 4.5 Isosurfaces of the thin-plate spline estimates
Figure 11: $\pm 0.3^{\circ}\mathrm{C}$ isosurfaces of the thin-plate spline fits, where negative temperature differences are shaded in blue and positive temperature differences are shaded in red. Each subplot is labeled with its camera perspective, given in its polar ($\theta$) and azimuthal ($\varphi$) angles.
For each set of isosurfaces, we first mask the thin-plate spline estimates that fail to reject a pointwise level $\alpha=0.05$ hypothesis test, using the variance estimates from Section 3.5.2. We then apply the marching cubes algorithm (Lorensen and Cline, 1987; Lewiner et al., 2012; Virtanen et al., 2020) to identify a positive and a negative isosurface of magnitude $0.3^{\circ}\mathrm{C}$. The positive isosurface is colored in red, and the negative in blue.
@@ 4.6 Vertically averaged temperature differences
The seasonally adjusted temperature differences and accompanying thin-plate spline fits are presented in Figure 12. In Figure 12, we only show data from pairs associated with hurricane-strength TCs, in contrast to Figure 6, which showed data from all profile pairs regardless of storm intensity. The restriction to hurricane-strength profile pairs here is to provide a fairer comparison to the thin-plate spline estimates of Figures 12 and 12, which use the hurricane-strength subset (as in Figure 7).
The thin-plate spline fit in Figure 12 conforms with our intuition; we observe a cooling signal, as in Figure 7, but its magnitude is dampened by the vertical averaging procedure. More importantly, Figure 12 does not show any warming effect indicative of vertical mixing. This is in contrast to Figures 7, 9, and 10, which treat depth as a separate dimension of analysis. From an oceanographic and meteorological perspective, this suggests that the vertical mixing-induced warming effect is local in depth; it only appears at some pressure levels, because it is caused by the forcing of warm surface waters downward by the TC. Once we integrate vertically, the local warming effect is cancelled out by a cooling effect near the surface and is no longer visible. From a statistical perspective, this underscores the importance of treating depth as a separate axis of analysis, using the vertical resolution of Argo observations.
@@
Aside from methodological contributions, we have presented several results of scientific interest. In particular, our framework improves upon past scientific literature by modeling the tropical cyclone-induced ocean thermal response as a function of time, pressure, and cross-track angle, in which both the time and cross-track angle components are treated continuously. Through this modeling, we are able to produce fine-grained characterizations of the ocean thermal response at a number of different depths, and we observe differing ocean thermal response as a function of depth. Specifically, we recover the primary effect of surface cooling and secondary effect of subsurface warming due to vertical mixing. By computing isosurfaces on our thin-plate spline fits, we are able to characterize these two phenomena as a function of the depth, time, and cross-track distance dimensions. Finally, the seasonal mean field component of our methodological framework reveals a nontrivial warming of the ocean near its surface during the TC season, which we estimate and account for in our analysis.
@@
The present methodological framework also admits a number of possible refinements. From a statistical standpoint, one could improve upon our model by considering each temperature profile as a functional observation and modeling the depth component continuously, rather than interpolating to a sequence of depths and fitting separate models at each depth level. This would potentially allow one to also infer changes in mixed layer depth and thermocline stability. Our model could also potentially be improved by replacing the thin-plate spline smoother with a spline smoother with a spatially varying smoothing parameter (e.g., Ruppert and Carroll, 2000) or a nonlinear, multiresolution smoother that naturally accounts for heterogeneity in the TC-induced signal (e.g., Daubechies et al., 1999).
@@
An additional extension of statistical and scientific interest involves carefully accounting for float drift. In this paper, we require the baseline and signal profiles to be within $0.2^{\circ}$ of each other and then assume that they share a single location in the analysis, admitting a useful simplification of the fitted covariance function. An extension would involve relaxing the strict closeness requirement of the paired profiles and removing the assumption that the paired profiles are in the same location by more carefully modeling their covariance, taking into account their relative spatial locations. This requires a nontrivial generalization of the Gaussian process model of Kuusela and Stein (2018), but would provide the benefit of increasing the number of available pairs with which to fit the model. An early approach in this vein is that of Paciorek and Schervish (2006). Yet another extension would be the consideration of parametric models for the TC wake (Haney et al., 2012). Finally, the fusion of microwave sea surface temperature (mSST; Gentemann et al., 2003; Wentz et al., 2005) observations with Argo profiles might provide an opportunity to produce estimates on the individual-storm level, rather than in terms of the global aggregates, as was done here. mSST observations are very dense in latitude and longitude, but lack the depth penetration provided by the Argo profiles, which are relatively sparse in latitude and longitude. Therefore, mSST and Argo are naturally complementary in the information provided.
@@
Appendix A Derivation of the thin-plate spline estimator
In this section, provided for completeness, we motivate and derive the thin-plate spline estimator (Duchon, 1977; Wahba, 1980, 1990; Green and Silverman, 1994; Wood, 2003) used in Section 3.5 for completeness. First, we review the natural thin-plate spline smoother in Section A.1. Sections A.2 and A.3 then describe versions of the thin-plate spline when subject to covariance reweighting (via generalized least squares; Rao, 1973; Draper and Smith, 1998; Nychka, 2000; Ruppert et al., 2003) and when its basis functions are constructed from a fixed knot set (Wahba, 1980; Wood, 2017).
A.1 The natural thin-plate spline smoother